OpenStudy (anonymous):

Factor completely: 2t²+5t-3 Please explain how you did it! Thanks!

OpenStudy (anonymous):

(2t -1 )( t+3 )

OpenStudy (anonymous):

2t^2 + 2t + 3t - 3 2t(t + 1) + -3(t+1) (2t -1)( t+3 )

OpenStudy (anonymous):

how did you get that? I need to show work

OpenStudy (anonymous):

Never mind

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

Actually no I'm not. Could you explain how you got that?

OpenStudy (anonymous):

sorry to be so long - i had to do a chore ok this is the way i do them what we need is two numbers which multiplied together with 2 = -6 and when added give + 5 make 3 columns and try out the combinations 2 -3 +5 aiming for +5 --------------------------------------- 2 x 1 3 x -1 2 x -1 + 1x3 = =1 2 x 1 3 x -1 3 x 2 + 1 x-1 = 6-1 = 5 - that is 3 with 2 and 1 with -1 so we get (2t - 1)(t + 3) outside 2t and 3 and inside -1 and 1t

OpenStudy (anonymous):

so there can be three numbers?

OpenStudy (anonymous):

I got confused after you put that the 2t is outside when its put inside,

OpenStudy (anonymous):

in this case there are 3 numbers involved the 2 from 2t^2, the -3 and the 5 from 5t. things are easier if the expression starts with x^2 then you only need two columns

OpenStudy (anonymous):

yeah I know about that! I actually get those ones!

OpenStudy (anonymous):

The ones with only x^2

OpenStudy (anonymous):

what i meant was the +5t is formed from the two 'outside' numbers 2t and +3 and the 'inside' ones -1 and t

OpenStudy (anonymous):

Alright, I kind of get it, the only thing is I have to show my work and I have no idea how to show this!

OpenStudy (anonymous):

well - just write the 3 columns on right side of your page - its a valid method and the one most favoured in the uk. i used to teach maths and i found most pupils understood this one best.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i find its better as you are writing down rather trying to figure out in your head - its not the only way to proceed of course - in end its a matter of choice, what you find easiest

OpenStudy (anonymous):

Yeah I definitely can't do stuff in my head by sides simple multiplication, adding subtracting, and dividing! Thanks so much for explaining!

OpenStudy (anonymous):

ok - practice them and you'll find you'll get good at it